Bridgeport Company purchased a delivery truck for $27,000 on January 1, 2020. The truck has an expected salvage value of $1,500, and is expected to be driven 102,000 miles over its estimated useful life of 10 years. Actual miles driven were 12,800 in 2020 and 14,300 in 2021.

'Calculate depreciation expense per mile under units-of-activity method. (Round answer to 2 decimal places, e.g. 0.50.)

Depreciation expense $

per mile
Answer #1

Depreciation expense = (Cost - Salvage value) /Total miles

= (27,000-1500)/102,000

= 25,500/102,000

= 0.25 per mile

2020 = 12,800*0.25 = 3200

2021 = 14,300*0.25 = 3575